### v1

@@ 3 Boundary integral operators
In this paper, we consider the scattering of time-harmonic electromagnetic fields with angular frequency $\omega>0$. The exterior region $\Omega^{c}$ is assumed to be filled by a homogeneous, isotropic and linear material with constant permittivity $\epsilon>0$ and permeability $\mu>0$. The wave number is defined by $\kappa:=\omega\sqrt{\epsilon\mu}>0$. In aid of further analysis with integral operators of a purely imaginary wave number, we state “the electric wave equation” and the Silver-Müller radiation condition with wave number $\sigma$, which can be either $\kappa$ or $i\kappa$,
@@
When $\sigma=\kappa$, the problem (2)-(3), in conjunction with the Dirichlet boundary condition
@@
describes the perfect conductor scattering at the boundary $\Gamma$ [14, Section 6.4]. In this case, the unknown $\boldsymbol{e}$ represents the scattered electric field, and the incident electric field $\boldsymbol{e}^{in}$ plays the role of excitation. The exterior problem has a unique solution $\boldsymbol{e}$ for all wave number $\kappa$ according to Rellich’s lemma [10, 29], whereas the interior problem is only uniquely solvable if $\kappa^{2}$ does not coincide with one of the Dirichlet or Neumann eigenvalues of the curlcurl-operator inside $\Omega$.
When $\sigma=i\kappa$, the equation (2) is called the Yukawa-type equation. More details on the properties of integral operators associated with this equation can be found in [29, Section 5.6.4], [21, Section 5.1] and [34].
@@
$G_{\sigma}(\boldsymbol{x},\boldsymbol{y}):=\dfrac{\exp(i\sigma\left|{\boldsymbol{x}-\boldsymbol{y}}\right|)}{4\pi\left|{\boldsymbol{x}-\boldsymbol{y}}\right|},$
for $\boldsymbol{x}\neq\boldsymbol{y}$. Then, the scalar and vectorial single layer potentials are respectively defined by
@@
with $\boldsymbol{x}\notin\Gamma$. In the electromagnetic scattering problem, we are more interested in the Maxwell single and double layer potentials
@@
which continuously map $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ into $\operatorname{\mathbf{H}}_{\textup{loc}}(\textbf{{curl}}^{2},\Omega\cup\Omega^{c})$, see [21, Theorem 17]. For any $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, $\boldsymbol{\Psi}^{\sigma}_{SL}(\boldsymbol{u})$ and $\boldsymbol{\Psi}^{\sigma}_{DL}(\boldsymbol{u})$ are solutions to the problem (2)-(3). Therefore, it immediately holds that
@@
On the other hand, any solution $\boldsymbol{u}\in\operatorname{\mathbf{H}}_{\textup{loc}}(\textbf{{curl}}^{2},\Omega\cup\Omega^{c})$ to the problem (2)-(3) satisfies the Stratton-Chu representation formula (see, e.g., [14, Theorem 6.2], [29, Theorem 5.5.1] and [21, Theorem 22])
(6) $\boldsymbol{u}(\boldsymbol{x})=-\boldsymbol{\Psi}^{\sigma}_{DL}(\left[\gamma_{D}\right]_{\Gamma}\boldsymbol{u})(\boldsymbol{x})-\boldsymbol{\Psi}^{\sigma}_{SL}(\left[\gamma_{N}\right]_{\Gamma}\boldsymbol{u})(\boldsymbol{x}),$
for $\boldsymbol{x}\in\Omega\cup\Omega^{c}$, where $\left[\gamma\right]_{\Gamma}:=\gamma^{+}-\gamma^{-}$ is the jump of some trace $\gamma$ across the boundary $\Gamma$, and the superscripts -- and $+$ designate traces onto $\Gamma$ from $\Omega$ and $\Omega^{c}$, respectively. In addition, the following jump relations hold true [21, Theorem 18]
(7) $\left[\gamma_{D}\right]_{\Gamma}\circ\boldsymbol{\Psi}^{\sigma}_{SL}=\boldsymbol{0},\qquad\qquad\qquad\left[\gamma_{N}\right]_{\Gamma}\circ\boldsymbol{\Psi}^{\sigma}_{SL}=-Id,$
@@
For $\sigma=\kappa$ or $\sigma=i\kappa$, with $\kappa>0$, we define
@@
By means of the relations (5), the tangential and Neumann traces of the Maxwell double layer potential $\boldsymbol{\Psi}^{\sigma}_{DL}$ can be derived accordingly. Then, the jump relations (7) allow us to deduce that
@@
Based on the regularity of the kernel $G_{\kappa}-G_{i\kappa}$, the following result can be obtained by following the same lines of the proof of [29, Theorem 3.4.1] or [9, Theorem 3.12]. This operator forms a compact perturbation of the integral operator $C_{\kappa}$ with principal part $C_{i\kappa}$.
@@
For $\kappa>0$, the integral operator $\delta C_{\kappa}:=C_{\kappa}-C_{i\kappa}$ is compact.
@@
Now, we examine the properties of the boundary integral operators with purely imaginary wave number $\sigma=i\kappa$, namely $S_{i\kappa}$ and $C_{i\kappa}$. The following lemma is taken from [21, Theorem 21] and [34, Theorem 2.5].
@@
Lemma 3.3 ($\operatorname{\mathbf{H}}_{\times}^{-1/2}(\textup{div}_{\Gamma},\Gamma)$-ellipticity of $S_{i\kappa}$).
@@
For $\kappa>0$, the integral operator $S_{i\kappa}$ is $\operatorname{\mathbf{H}}_{\times}^{-1/2}(\textup{div}_{\Gamma},\Gamma)$-elliptic and symmetric with respect to ,×$\langle{\cdot,\cdot}\rangle_{\times,\Gamma}$, i.e., there exists a constant $C>0$ depending only on $\Gamma$ and $\kappa$ such that, for all $\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$
@@
$\left\langle{\boldsymbol{v},S_{i\kappa}\boldsymbol{u}}\right\rangle_{\times,\Gamma}=\left\langle{\boldsymbol{u},S_{i\kappa}\boldsymbol{v}}\right\rangle_{\times,\Gamma},\qquad\left\langle{\boldsymbol{u},S_{i\kappa}\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}\geq C\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)}.$
The ellipticity of $S_{i\kappa}$ implies that its inverse $S_{i\kappa}^{-1}:\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)\to\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ exists and is continuous. Throughout the paper, the notations $C$ and $\widetilde{C}$ (without subscripts) are used for generic positive constants which depend only on $\Gamma$ and $\kappa$. In general, these notations can have different values in different contexts. The next result is crucial in proving the coercivity of the MFIE operator in the next section.
Lemma 3.4 ($S_{i\kappa}$-coercivity).
For $\kappa>0$, there exists a constant $C_{A}>0$ depending only on $\Gamma$ and $\kappa$ such that, for all $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$
@@
(8) $\left\langle{\left({\dfrac{1}{2}Id+C_{i\kappa}}\right)\boldsymbol{u},S_{i\kappa}\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}\geq C_{A}\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)}.$
@@
The integration by parts formula (1) and the relations (5) allow us to deduce that
$\left\langle{\left({\dfrac{1}{2}Id+C_{i\kappa}}\right)\boldsymbol{u},S_{i\kappa}\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}$ $=\left\langle{\left({\gamma_{D}^{-}\circ\textbf{{curl}}\circ\boldsymbol{\Psi}^{i\kappa}_{SL}}\right)(\boldsymbol{u}),\left({\gamma_{D}^{-}\circ\boldsymbol{\Psi}^{i\kappa}_{SL}}\right)(\overline{\boldsymbol{u}})}\right\rangle_{\times,\Gamma}$
$=\int_{\Omega}\left({\left|{\textbf{{curl}}\,\boldsymbol{\Psi}^{i\kappa}_{SL}(\boldsymbol{u})}\right|^{2}+\kappa^{2}\left|{\boldsymbol{\Psi}^{i\kappa}_{SL}(\boldsymbol{u})}\right|^{2}}\right)\,\mathrm{d}\boldsymbol{x}$
$\geq\min\{1,\kappa^{2}\}\left\lVert{\boldsymbol{\Psi}^{i\kappa}_{SL}(\boldsymbol{u})}\right\rVert^{2}_{\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)}$
@@
$\geq C\left\lVert{S_{i\kappa}(\boldsymbol{u})}\right\rVert^{2}_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)}$
@@
$\geq C_{A}\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)}.$
In the last two inequalities, we used the continuity of the tangential trace $\gamma_{D}^{-}$ and the operator $S_{i\kappa}^{-1}$. Hence, we complete the proof.

### v2

@@ 3 Boundary integral operators
In this paper, we consider the scattering of time-harmonic electromagnetic fields with angular frequency $\omega>0$. The exterior region $\Omega^{c}$ is assumed to be filled by a homogeneous, isotropic and linear material with constant permittivity $\epsilon>0$ and permeability $\mu>0$. The wave number is defined by $\kappa:=\omega\sqrt{\epsilon\mu}>0$. In aid of further analysis with integral operators of a purely imaginary wave number, we start from the “the electric wave equation” and the Silver-Müller radiation condition with generic wave number $\sigma$, which can be either $\kappa$ or $i\kappa^{\prime}$, with $\kappa,\kappa^{\prime}>0$
@@
When $\sigma=\kappa$, the problem (2)-(3) in conjunction with the Dirichlet boundary condition
@@
describes the scattering by a perfect electric conductor at the boundary $\Gamma$ [16, Section 6.4]. In this case, the unknown $\boldsymbol{e}$ represents the scattered electric field, and the incident electric field $\boldsymbol{e}^{in}$ plays the role of excitation. The exterior problem has a unique solution $\boldsymbol{e}$ for all wave number $\kappa$ according to Rellich’s lemma [11, 33], whereas the interior problem is only uniquely solvable if $\kappa^{2}$ does not coincide with one of the Dirichlet or Neumann eigenvalues of the curlcurl-operator inside $\Omega$.
When $\sigma=i\kappa^{\prime}$, the equation (2) is called the Yukawa-type equation. More details on the properties of boundary integral operators associated with this equation can be found in [33, Section 5.6.4], [24, Section 5.1] and [39].
@@
$G_{\sigma}(\boldsymbol{x},\boldsymbol{y}):=\dfrac{\exp(i\sigma\left|{\boldsymbol{x}-\boldsymbol{y}}\right|)}{4\pi\left|{\boldsymbol{x}-\boldsymbol{y}}\right|},\qquad\qquad\boldsymbol{x}\neq\boldsymbol{y}.$
Then, the scalar and vectorial single layer potentials are respectively defined by
@@
with $\boldsymbol{x}\notin\Gamma$. In electromagnetics, we are more interested in the Maxwell single and double layer potentials
@@
which continuously map $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ into $\operatorname{\mathbf{H}}_{\textup{loc}}(\textbf{{curl}}^{2},\Omega\cup\Omega^{c})$, see [24, Theorem 17]. For any $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, the potentials $\boldsymbol{\Psi}^{\sigma}_{SL}(\boldsymbol{u})$ and $\boldsymbol{\Psi}^{\sigma}_{DL}(\boldsymbol{u})$ are solutions to the problem (2)-(3). Therefore, it immediately holds that
@@
On the other hand, every solution $\boldsymbol{u}\in\operatorname{\mathbf{H}}_{\textup{loc}}(\textbf{{curl}}^{2},\Omega\cup\Omega^{c})$ to the problem (2)-(3) satisfies the Stratton-Chu representation formula (see, e.g., [16, Theorem 6.2], [33, Theorem 5.5.1] and [24, Theorem 22])
(6) $\boldsymbol{u}(\boldsymbol{x})=-\boldsymbol{\Psi}^{\sigma}_{DL}(\left[\gamma_{D}\right]_{\Gamma}\boldsymbol{u})(\boldsymbol{x})-\boldsymbol{\Psi}^{\sigma}_{SL}(\left[\gamma_{N}\right]_{\Gamma}\boldsymbol{u})(\boldsymbol{x}),\qquad\quad\boldsymbol{x}\in\Omega\cup\Omega^{c},$
where $\left[\gamma\right]_{\Gamma}:=\gamma^{+}-\gamma^{-}$ is the jump of some trace $\gamma$ across the boundary $\Gamma$, and the superscripts -- and $+$ designate traces onto $\Gamma$ from $\Omega$ and $\Omega^{c}$, respectively. In addition, the following jump relations hold true [24, Theorem 18]
(7) $\left[\gamma_{D}\right]_{\Gamma}\circ\boldsymbol{\Psi}^{\sigma}_{SL}=0,\qquad\qquad\qquad\left[\gamma_{N}\right]_{\Gamma}\circ\boldsymbol{\Psi}^{\sigma}_{SL}=-Id,$
@@
For $\sigma=\kappa$ or $\sigma=i\kappa^{\prime}$, with $\kappa,\kappa^{\prime}>0$, we define
@@
By means of the relations (5), the tangential and Neumann traces of the Maxwell double layer potential $\boldsymbol{\Psi}^{\sigma}_{DL}$ can be derived accordingly. Then, the jump relations (7) imply that
@@
The following compactness is essential for further analysis of the MFIE, as it allows us to primarily work with the operator $C_{i\kappa^{\prime}}$ and then transfer to $C_{\kappa}$ by a compact perturbation argument. This result can be obtained by following the approach in the proof of [33, Theorem 3.4.1] or [10, Theorem 3.12], based on the regularity of the kernel $G_{\kappa}-G_{i\kappa^{\prime}}$.
@@
For any $\kappa,\kappa^{\prime}>0$, the operator $C_{\kappa}-C_{i\kappa^{\prime}}$ is compact.
@@
Now, we examine the properties of the boundary integral operators with purely imaginary wave number $\sigma=i\kappa^{\prime}$, namely $S_{i\kappa^{\prime}}$ and $C_{i\kappa^{\prime}}$. The following lemma is taken from [24, Theorem 21] and [39, Theorem 2.5].
@@
Lemma 3.3 ($\operatorname{\mathbf{H}}_{\times}^{-1/2}(\textup{div}_{\Gamma},\Gamma)$-ellipticity of $S_{i\kappa^{\prime}}$).
@@
For $\kappa^{\prime}>0$, the integral operator $S_{i\kappa^{\prime}}$ is symmetric with respect to ,×$\langle{\cdot,\cdot}\rangle_{\times,\Gamma}$ and $\operatorname{\mathbf{H}}_{\times}^{-1/2}(\textup{div}_{\Gamma},\Gamma)$-elliptic, i.e., there exists a constant $C>0$ depending only on $\kappa^{\prime}$ and $\Gamma$ such that
@@
$\left\langle{\boldsymbol{v},S_{i\kappa^{\prime}}\boldsymbol{u}}\right\rangle_{\times,\Gamma}=\left\langle{\boldsymbol{u},S_{i\kappa^{\prime}}\boldsymbol{v}}\right\rangle_{\times,\Gamma},$ $\forall\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma),$
@@
$\left\langle{\boldsymbol{u},S_{i\kappa^{\prime}}\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}\geq C\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)},$ $\forall\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma).$
The symmetry and ellipticity of $S_{i\kappa^{\prime}}$ imply that the space $\operatorname{\mathbf{X}}_{\kappa^{\prime}}:=\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ equipped with the inner product
$\left({\boldsymbol{u},\boldsymbol{v}}\right)_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}:=\left\langle{\boldsymbol{u},S_{i\kappa^{\prime}}\overline{\boldsymbol{v}}}\right\rangle_{\times,\Gamma},\qquad\qquad\quad\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma),$
is a Hilbert space. Its induced norm is denoted by $\left\lVert{\cdot}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$. Please note that the two spaces $\operatorname{\mathbf{X}}_{\kappa^{\prime}}$ and $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ only differ in their equipped norms. In what follows, the space $\operatorname{\mathbf{X}}_{\kappa^{\prime}}$ is only used when its inner product and norm are involved or for a compact notation, whereas $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ is used as the space of their elements (independent of $\kappa^{\prime}$).
The next results concern the contraction properties of the operators $2C_{i\kappa^{\prime}}$ and $\tfrac{1}{2}Id\pm C_{i\kappa^{\prime}}$. The latter were shown in [39] by extending the idea in [38, Theorem 3.1] for Laplace’s equation. In the following, we prove the contraction of $2C_{i\kappa^{\prime}}$ with respect to the norm $\left\lVert{\cdot}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$ by means of the Calderón projection formulas for Maxwell’s equations.
Lemma 3.4 (Contraction property).
For any $\kappa^{\prime}>0$, there holds that
(8) $\left\langle{\boldsymbol{v},C_{i\kappa^{\prime}}\boldsymbol{u}}\right\rangle_{\times,\Gamma}=\left\langle{\boldsymbol{u},C_{i\kappa^{\prime}}\boldsymbol{v}}\right\rangle_{\times,\Gamma},$ $\forall\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma),$
(9) $\left\lVert{C_{i\kappa^{\prime}}\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\leq\beta\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}},$ $\forall\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma),$
where the constant $\beta>0$ is defined by
(10) $\beta(\kappa^{\prime},\Gamma):=\sqrt{\dfrac{1}{4}-\dfrac{{\kappa^{\prime}}^{2}}{C_{S}^{2}}}<\dfrac{1}{2},\qquad\quad C_{S}(\kappa^{\prime},\Gamma):=\sup_{\boldsymbol{u}\in\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\dfrac{\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}{\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}.$
@@
The symmetry of $C_{i\kappa^{\prime}}$ with respect to ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$ can be obtained by following the lines in the proof of [10, Theorem 3.9], with the help of the integration by parts formula (1). For the contraction property of $2C_{i\kappa^{\prime}}$, we need the following Calderón projection formulas, which are derived from the Stratton-Chu representation formula (6) for $\sigma=i\kappa^{\prime}$ and [8, Formulation 35]
$C_{i\kappa^{\prime}}\circ S_{i\kappa^{\prime}}$ $=-S_{i\kappa^{\prime}}\circ C_{i\kappa^{\prime}},$
$C_{i\kappa^{\prime}}\circ C_{i\kappa^{\prime}}$ $=\dfrac{1}{4}Id+{\kappa^{\prime}}^{2}S_{i\kappa^{\prime}}\circ S_{i\kappa^{\prime}}.$
Based on these identities and the symmetry of $S_{i\kappa^{\prime}}$ and $C_{i\kappa^{\prime}}$, we can deduce that
@@
$\left\lVert{C_{i\kappa^{\prime}}\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$ $=\left\langle{C_{i\kappa^{\prime}}\boldsymbol{u},\left({S_{i\kappa^{\prime}}\circ C_{i\kappa^{\prime}}}\right)\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}$
$=\left\langle{\boldsymbol{u},\left({C_{i\kappa^{\prime}}\circ C_{i\kappa^{\prime}}\circ S_{i\kappa^{\prime}}}\right)\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}$
$=\dfrac{1}{4}\left\langle{\boldsymbol{u},S_{i\kappa^{\prime}}\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}+{\kappa^{\prime}}^{2}\left\langle{\boldsymbol{u},\left({S_{i\kappa^{\prime}}\circ S_{i\kappa^{\prime}}\circ S_{i\kappa^{\prime}}}\right)\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}$
@@
$=\dfrac{1}{4}\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}-{\kappa^{\prime}}^{2}\left\lVert{S_{i\kappa^{\prime}}\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$
$\leq\left({\dfrac{1}{4}-\dfrac{{\kappa^{\prime}}^{2}}{C_{S}^{2}}}\right)\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}.$
The contraction properties of $\tfrac{1}{2}Id\pm C_{i\kappa^{\prime}}$ with respect to $\left\lVert{\cdot}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$ can be achieved as a consequence of (9).
Corollary 3.6.
For any $\kappa^{\prime}>0$, the operators $\tfrac{1}{2}Id\pm C_{i\kappa^{\prime}}$ are contractions with respect to the norm $\left\lVert{\cdot}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$. In particular, for all $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, it holds that
(11) $\left({\dfrac{1}{2}-\beta}\right)\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\leq\left\lVert{\left({\dfrac{1}{2}Id\pm C_{i\kappa^{\prime}}}\right)\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\leq\left({\dfrac{1}{2}+\beta}\right)\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}.$
The following $S_{i\kappa^{\prime}}$-coercivity of $\tfrac{1}{2}Id\pm C_{i\kappa^{\prime}}$ is crucial in proving the coercivity of the MFIE operator in the next section.
Lemma 3.7 ($S_{i\kappa^{\prime}}$-coercivity).
For any $\kappa^{\prime}>0$, there holds that
(12) $\left({\left({\dfrac{1}{2}Id\pm C_{i\kappa^{\prime}}}\right)\boldsymbol{u},\boldsymbol{u}}\right)_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\geq\left({\dfrac{1}{4}-\beta^{2}}\right)\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}},\qquad\forall\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma).$
Proof 3.8.
For all $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, it holds that
$\left({\left({\dfrac{1}{2}Id\pm C_{i\kappa^{\prime}}}\right)\boldsymbol{u},\boldsymbol{u}}\right)_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$ $=\dfrac{1}{2}\left({\boldsymbol{u},\boldsymbol{u}}\right)_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\pm\left({C_{i\kappa^{\prime}}\boldsymbol{u},\boldsymbol{u}}\right)_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$
$\geq\dfrac{1}{2}\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}-\left\lVert{C_{i\kappa^{\prime}}\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{\boldsymbol{u}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$
$\geq\dfrac{1}{2}\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}-\dfrac{1}{4}\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}-\left\lVert{C_{i\kappa^{\prime}}\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$
$\geq\left({\dfrac{1}{4}-\beta^{2}}\right)\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}.$
Remark 3.9.
The contraction properties (9) and (11) remain valid when replacing $\left\lVert{\cdot}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$ by the equivalent norm $\left\lVert{S^{-1}_{i\kappa^{\prime}}\cdot}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$. The space $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ equipped with this norm also defines a Hilbert space. In addition, analogous to (12), the operators $\tfrac{1}{2}Id\pm C_{i\kappa^{\prime}}$ are also $S_{i\kappa^{\prime}}^{-1}$-coercive, i.e., for all $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, it holds that
$\left\langle{\left({S^{-1}_{i\kappa^{\prime}}\circ\left({\dfrac{1}{2}Id\pm C_{i\kappa^{\prime}}}\right)}\right)\boldsymbol{u},\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}\geq\left({\dfrac{1}{4}-\beta^{2}}\right)\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{u}}\right\rVert^{2}_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}.$